OpenStudy (anonymous):

In which polygon is the sum of the measures of the interior angles 2160°? A. Nonagon B. 12-gon C. Decagon D. 14-gon

OpenStudy (anonymous):

The sum of the interior angles of any polygon of n-sides is: 180(n - 2)

OpenStudy (anonymous):

So it is 180(14-2)

OpenStudy (anonymous):

2160 = 180(n - 2) Divide both sides by 180, you get 12 = n -2 n = 14

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Welcome.